Julie Convenience Store sold merchandise for cash to a customer, and recorded a debit to Cash for $371, which included a 6% Sales tax. In the same transaction, they must also:

A) credit Sales Revenue for 300

B) credit Sales Tax Payable for $22.26

C) credit Sales Tax Payable for $21

D) credit Sales Revenue for $371

E) credit Sales Revenue for $393.26
Answer #1
Answer: Option [C] Credit sales tax payable for $21.
Calculation and explanation:
The sales tax element = 371*6/106 = $21
The price exclusive of sales tax would be 371-21 = $350
The correct entry should be:
Cash                                       371
Sales revenue    350
Sales tax payable 21